OpenStudy (shawnnelson):

Suppose that ln2=a and ln3=b. Use properties of logarithms to write each log in terms of a and b. -> ln6

jhonyy9 (jhonyy9):

6=3*2 yes ?

jhonyy9 (jhonyy9):

do you know this property of logarithms ? log a +log b = log a*b

OpenStudy (shawnnelson):

I am not sure to be honest but i have the book with me and it shows me what it should be but i just want to know how to work it out. The actual answer is a+b

jhonyy9 (jhonyy9):

ln6 = ln(3*2) ln(3*2) = ln3 +ln2 ok. ?

jhonyy9 (jhonyy9):

ln6 = ln3 +ln2

OpenStudy (shawnnelson):

ok and waht if it was b-a